-10 PM Mon Jan 21 Done Ch. 1- Need-to-Know Videos Help Save & Ext Submit QS 1-10 ldentifying effects of transactions using accounting equation-Revenues and Expenses LO P1 10 The following transactions were completed by the company. a. The company completed consulting work for a client and immediately collected $5,500 cash eamed b. The company completed commission work for a client and sent a bill for $4,000 to be received within 30 days c. The company paid an assistant $1,400 cash as wages for the period d. The company collected $1,000 cash as a partial payment for the amount owed by the client in transaction b e. The company paid $700 cash for this periods cleaning services. Required: Enter the impact of each transaction on individual items of the accounting equation. (Enter decreases to account balances with minus sign) ferences Cash Accounts Common Bal <Prev 8of10 Next >
Answer #1

Impact of each transaction on individual items of the accounting equation is as shown below:

Assets = Liabilities + Equity
Event Number Cash + Account Receivable = Accounts Payable Common Stock - Dividends + Revenue - Expense
a 5,500 = 5,500 -
b 4,000 = 4,000 -
Bal. 5,500 4,000 = 9,500 -
c. -1,400 = - 1,400
Bal. 4,100 4,000 = 9,500 - 1,400
d 1,000 -1,000 = -
Bal. 5,100 3,000 = 9,500 - 1,400
e. -700 = - 700
Bal. 4,400 3,000 = 9,500 - 2,100